# Combinatorial Creativity: A New Frontier in Generalization Abilities

### v4

@@ Abstract
Artificial intelligence (AI) systems, and Large Language Models (LLMs) in particular, are increasingly employed for creative tasks like scientific idea generation, constituting a form of generalization from training data unaddressed by existing conceptual frameworks. Despite its similarities to compositional generalization (CG), combinatorial creativity (CC) is an open-ended ability. Instead of evaluating for accuracy or correctness against fixed targets, which would contradict the open-ended nature of CC, we propose a theoretical framework and algorithmic task for evaluating outputs by their degrees of novelty and utility. From here, we make several important empirical contributions: (1) We obtain the first insights into the scaling behavior of creativity for LLMs. (2) We discover that, for fixed compute budgets, there exist optimal model depths and widths for creative ability. (3) We find that the ideation-execution gap, whereby LLMs excel at generating novel scientific ideas but struggle to ensure their practical feasibility, may be explained by a more fundamental novelty-utility tradeoff characteristic of creativity algorithms in general. Importantly, this tradeoff remains persistent even at scale, casting doubt on the long-term creative potential of LLMs in their current form. Together, our conceptual framework and empirical findings provide a foundation for understanding and improving creativity in modern AI models, bridging the gap between human and machine intelligence.
[…]
@@ 1 Introduction
To address these limitations, we introduce a formal framework and an open-ended, algorithmic task for evaluating combinatorial creativity. Our framework models creativity within a conceptual space, represented as a large graph where models must find novel paths between concepts while adhering to logical constraints. Using this testbed, we conduct a large-scale empirical study of decoder-only Transformers, which have become the de facto architecture for LLMs, by systematically varying their size, depth, and width to yield several key insights.
First, we obtain the first insights into the scaling behavior for creativity, demonstrating a predictable improvement in performance with increased model size and training compute. Second, we uncover a crucial architectural insight: for a fixed computational budget, there is an optimal model depth and width that maximizes creativity, beyond which performance diminishes. Third, we perform a detailed error analysis, which reveals that as task complexity increases, models are more likely to fail by violating utility constraints rather than by producing non-novel outputs. This is explained by our fourth and most critical finding: a fundamental novelty-utility tradeoff, which persists even as model scale increases. This suggests the ideation-execution gap may be an intrinsic challenge for current architectures—and not merely a problem of scale. Together, our conceptual framework and empirical findings offer a launchpad for understanding and improving the creativity of modern AI models, as we enter a new frontier in generalization abilities.
@@ 2.2 Distinguishing Combinatorial Creativity from Classical Forms of Generalization
Table 1: Comparison of forms of compositional generalization, productivity (CG-P) and systematicity (CG-S), with combinatorial creativity (CC) along six key dimensions. (A1) Compositionality: all three abilities always construct compositional objects; (A2) Open-Ended: $C$ is the only ability which must always be evaluated in an open-ended way, meaning there are always many ways to adequately solve a particular task; (A3) Structural Novelty: CG-P always involves generalizing to unseen lengths and structures, whereas this is only true of CG-S and $C$ sometimes; (A4) Semantic Novelty: CG-S and $C$ always involve combining primitives in a way that leads to semantically novel structures, whereas this is only true of CG-P sometimes; (A5) Degree of Novelty and (A6) Degree of Utility: $C$ is the only ability which always quantifies the novelty and utility of its artifacts in degrees, rather than by binary evaluation. On the right, we compare our framework in Section 3 against sibling discovery (SD) and triangle discovery (TD) from Nagarajan et al. (2025). A more detailed comparison of our framework and SD/TD is given in Section A.1.
@@ 3.1 Combinatorial Creativity Setting
A creative prompt is a tuple $x=(u,v,\mathcal{I},\mathcal{X})$ consisting of (i) a starting concept $u\in\mathcal{V}$, (ii) an ending concept $v\in\mathcal{V}$, (iii) an inclusion set $\mathcal{I}\subseteq\Sigma$, and (iv) an exclusion set $\mathcal{X}\subseteq\Sigma$ such that $\mathcal{I}\cap\mathcal{X}=\emptyset$.
@@ 3.2 Quantifying Degrees of Novelty
Given a creative artifact $P$ of length $h$, defined according to Definition 2, its novelty is given by:
@@
where $S(P)=\frac{1}{k}\sum_{i=1}^{k}-\log(w_{l_{i}})$ is the surprise of the path, defined as the average negative log-likelihood of the label probabilities given in Definition 1, and $\alpha_{h},\alpha_{r}>0$ are controllable, scalar parameters.
@@ 3.3 Quantifying Degrees of Utility
$\text{U}(P;x):=\left(1+\alpha_{I}|I|\right)\left(1+\alpha_{X}|X|\right)\mathbb{I}~\![v_{0}=u,v_{h}=v,\{\ell_{1},...,\ell_{h}\}\supseteq I,\{\ell_{1},...,\ell_{h}\}\cap X=\emptyset]$ (3)
@@
The utility function consists of three main parts: the terms $\left(1+\alpha_{I}|\mathcal{I}|\right)$ and $\left(1+\alpha_{X}|\mathcal{X}|\right)$ scale the utility function in proportion to the number of inclusion and exclusion constraints, respectively, while the indicator term ensures that artifacts obey these constraints and start and end at the correct nodes. The number of total constraints also allows us to informally define “utility levels:” level $1$ refers to unconstrained query ($I=\emptyset,X=\emptyset$), while levels $l>1$ each consist of l1l-1 constraints. In Section 5, we use these utility levels to demonstrate the novelty-utility tradeoff in creative LLM outputs. We also explain how the constraints can serve as minimal abstractions of failure modes encountered when using LLMs for scientifically creative idea generation.
@@ 3.4 Measuring Creativity
With this framework in place, we now investigate how creativity varies with architectural design and compute scaling.
@@ 4 Experiments
We are interested in how fundamental architectural choices influence the creativity of LLMs on the task defined in Section 3. For example, Nagarajan et al. (2025) recently found creative gains from changing the pre-training objective from next-token to multi-token prediction. In this study, we are especially curious whether creative limits present in smaller models also persist at scale, as this can provide insight into whether continued scaling of LLMs can yield even greater gains in creativity, or if architectural innovations are needed to support more open-ended behaviors (Hughes et al., 2024).
@@ 5 Results and Discussion
In Figure 3(a), we visualize the impact of the number of layers $L$ on combinatorial creativity across all three model sizes. Our most significant finding is that for a fixed parameter count, there is an architectural “sweet spot,” an optimal number of layers that maximizes creativity, after which increasing depth further can be detrimental. For the 100M models, this peak is clearly visible around 8 layers. Models that are too shallow (e.g., 2-4 layers) or too deep (e.g., 12+ layers) for their parameter count are substantially less creative. Similarly, in Figure 3(b), we visualize the impact of the width-to-depth ratio on the creativity of models at all three scales. Note that when depth is increased within a fixed parameter budget, the model’s width (embedding dimension) must necessarily decrease. For a fixed parameter count, there is also an optimal width-to-depth ratio that maximizes creativity, after which increasing the width further can be detrimental. The optimal $E/L$ ratio occurs between 200 and 300 for all three model sizes. This suggests that combinatorial creativity requires a delicate balance between (1) models that are too shallow and wide, where insufficient depth may hinder the sequential processing capacity to handle in-memory leaps of thought (which are required to make distant, constrained associations between concepts) and (2) models that are too deep and narrow, which suffer from restricted representational capacity that may limit their ability to hold and associate the diverse concepts needed for novel combinations.
@@
In Figure 4, we plot the relationship between novelty and utility across all three model sizes. Previously, Varshney (2019) established a fundamental, information-theoretic limit between novelty and utility for combinatorial creativity. We find a similar novelty-utility tradeoff holds here: across all three scales, as the number of utility constraints increases, the novelty of artifacts exhibits a clear downward trend. Importantly, this tradeoff does not improve by increasing model size, suggesting it is a fundamental characteristic of today’s LLMs.
@@
With the novelty-utility tradeoff in mind, we provide a potential explanation for why LLMs excel at generating novel research ideas (Si et al., 2024, Sanyal et al., 2025, Gu and Krenn, 2024, Wang et al., 2024, Radensky et al., 2024, Guo et al., 2025) but struggle at ensuring their practical feasibility, in what has been termed the ideation-execution gap (Si et al., 2025). In Table 2, we explain how our inclusion-exclusion constraints can serve as minimal abstractions for the key failure modes contributing to this gap. In particular, exclusion constraints can be viewed as a minimal abstraction for preventing unrealistic assumptions and excluding prohibitively expensive execution plans, while inclusion constraints can represent ensuring that a proper baseline is included and can serve as a minimal abstraction to ensure implementation plans are sufficiently detailed. Since the novelty-utility tradeoff remains persistent even at the 100M scale (see Figure 4), this suggests that the same fundamental tradeoff might also plague the frontier models used in previous works. This finding is consistent with recent work from Shashidhar et al. (2025), which also identified a validity-diversity tradeoff in LLM-generated evaluation questions, where models that produced the most diverse (novel) questions often did so at the cost of lower factual validity (utility).
@@ A.1 Open-Ended Algorithmic Tasks
Detailed Comparison with Sibling and Triangle Discovery
  1. 1.
  2. 2.
  3. 3.
@@ B.1 Dataset Construction
Evaluation Set Generation
For each base path found, we generate a hierarchy of $L_{\max}=6$ evaluation instances, or ”levels.”

### v5

@@ Abstract
Artificial intelligence (AI) systems, and Large Language Models (LLMs) in particular, are increasingly employed for creative tasks like scientific idea generation, constituting a form of generalization from training data unaddressed by existing conceptual frameworks. Despite its similarities to compositional generalization (CG), combinatorial creativity (CC) is an open-ended ability. Instead of evaluating for accuracy or correctness against fixed targets, which would contradict the open-ended nature of CC, we propose a theoretical framework and algorithmic task for evaluating outputs by their degrees of novelty and utility. From here, we make several important empirical contributions: (1) We obtain the first insights into the scaling behavior of creativity for LLMs. (2) We discover that, for fixed compute budgets, there exist optimal model depths and widths for creative ability. (3) We find that the ideation-execution gap, whereby LLMs excel at generating novel scientific ideas but struggle to ensure their practical feasibility, may be explained by a more fundamental novelty-utility tradeoff characteristic of creativity algorithms in general. Though our findings persist up to the 100M scale, frontier models today are well into the billions of parameters. Therefore, our conceptual framework and empirical findings can best serve as a starting point for understanding and improving the creativity of frontier-size models today, as we begin to bridge the gap between human and machine intelligence.
@@ 1 Introduction
To address these limitations in a controlled way, we introduce a formal framework and an open-ended, algorithmic task for evaluating combinatorial creativity. Our framework models creativity within a conceptual space represented as a large synthetic graph, where models must find novel paths between concepts while adhering to logical constraints. We use this as a minimal testbed that isolates structural aspects of creative generalization. Within this setting, we conduct a systematic empirical study of decoder-only Transformers, varying their size, depth, and width across 1M–100M parameters and training compute budgets to probe how these choices relate to creative performance.
First, we obtain initial evidence about the scaling behavior of combinatorial creativity, observing predictable improvements in performance with increased model size and training compute within our parameter regime. Second, we uncover an architectural trend: for a fixed computational budget on this task, wider, shallower models outperform deeper, narrower ones, with an intermediate depth–width tradeoff that maximizes creativity. Third, we perform a detailed error analysis, which reveals that as task complexity increases, models more often fail by violating utility constraints than by producing trivially non-novel outputs. Finally, we empirically recover a fundamental novelty–utility tradeoff predicted by prior theory (Varshney, 2019); in our experiments this tradeoff remains pronounced across all model sizes studied. These results do not aim to characterize the creative limits of frontier models but instead provide a controlled, algorithmic instance of phenomena—such as the tension between novelty and feasibility—that have been observed in scientific ideation with LLMs. Together, our conceptual framework and empirical findings offer a starting point for studying and improving the creativity of modern AI models, and for extending this line of work to larger scales and more semantically grounded conceptual spaces.
@@ 2.2 Distinguishing Combinatorial Creativity from Classical Forms of Generalization
Table 1: Comparison of forms of compositional generalization, productivity (CG-P) and systematicity (CG-S), with combinatorial creativity (CC) along six key dimensions. (A1) Compositionality: all three abilities always construct compositional objects; (A2) Open-Ended: $C$ is the only ability which must always be evaluated in an open-ended way, meaning there are always many ways to adequately solve a particular task; (A3) Structural Novelty: CG-P always involves generalizing to unseen lengths and structures, whereas this is only true of CG-S and $C$ sometimes; (A4) Semantic Novelty: CG-S and $C$ always involve combining primitives in a way that leads to semantically novel structures, whereas this is only true of CG-P sometimes; (A5) Degree of Novelty and (A6) Degree of Utility: $C$ is the only ability which always quantifies the novelty and utility of its artifacts in degrees, rather than by binary evaluation. On the right, we compare our framework in Section 3 against sibling discovery (SD) and triangle discovery (TD) from Nagarajan et al. (2025). A more detailed comparison of our framework and SD/TD is given in Section 3.5.
@@ 3.1 Combinatorial Creativity Setting
A creative prompt is a tuple $x=(u,v,\mathcal{I},\mathcal{X})$ consisting of (i) a starting concept $u\in\mathcal{V}$, (ii) an ending concept $v\in\mathcal{V}$, (iii) an inclusion set $\mathcal{I}\subseteq\Sigma$ of edges that must be present in the path, and (iv) an exclusion set $\mathcal{X}\subseteq\Sigma$ of edges that must be excluded from the path, such that $\mathcal{I}\cap\mathcal{X}=\emptyset$.
@@ 3.2 Quantifying Degrees of Novelty
Given a non-uniform distribution over edge labels $\mathbf{w}\in\Delta\Sigma$ and a creative artifact $P$ of length $h$, defined according to Definition 2, its novelty is given by:
@@
where $S(P)=\frac{1}{k}\sum_{i=1}^{k}-\log(w_{l_{i}})$ is the surprise of the path, defined as the average negative log-likelihood of the label probabilities $w_{l_{i}}$ given in Definition 1, and $\alpha_{h},\alpha_{r}>0$ are controllable, scalar parameters.
@@ 3.3 Quantifying Degrees of Utility
$\text{U}(P;x):=\left(1+\alpha_{I}|I|\right)\left(1+\alpha_{X}|X|\right)\mathbb{I}\penalty 10000\ \![v_{0}=u,v_{h}=v,\{\ell_{1},...,\ell_{h}\}\supseteq I,\{\ell_{1},...,\ell_{h}\}\cap X=\emptyset]$ (3)
@@
The utility function consists of three main parts: the terms $\left(1+\alpha_{I}|\mathcal{I}|\right)$ and $\left(1+\alpha_{X}|\mathcal{X}|\right)$ scale the utility function in proportion to the number of inclusion and exclusion constraints, respectively, while the indicator term ensures that artifacts obey these constraints and start and end at the correct nodes.
Evaluation Set Generation
For each base path found, we generate a hierarchy of $L_{\max}=5$ evaluation instances, or ”levels.”
@@ 3.4 Measuring Creativity
3.5 Detailed Comparison with Sibling and Triangle Discovery
  1. 1.
  2. 2.
  3. 3.
@@ 4 Experiments
We are interested in how fundamental architectural choices influence the creativity of LLMs on the task defined in Section 3. For example, Nagarajan et al. (2025) recently found creative gains from changing the pre-training objective from next-token to multi-token prediction. In this study, we are especially curious how model creativity is impacted by scale and architecture choice
@@ 5 Results and Discussion
In Figure 3(a), we visualize the impact of the number of layers $L$ on combinatorial creativity across all three model sizes. Our most significant finding is that for a fixed parameter count, there is an architectural “sweet spot,” an optimal number of layers that maximizes creativity, after which increasing depth further can be detrimental. For the 100M models, this peak is clearly visible around 8 layers. Models that are too shallow (e.g., 2-4 layers) or too deep (e.g., 12+ layers) for their parameter count are substantially less creative. Similarly, in Figure 3(b), we visualize the impact of the width-to-depth ratio on the creativity of models at all three scales. Note that when depth is increased within a fixed parameter budget, the model’s width (embedding dimension) must necessarily decrease. For a fixed parameter count, there is also an optimal width-to-depth ratio that maximizes creativity, after which increasing the width further can be detrimental. The optimal $E/L$ ratio occurs between 200 and 300 for all three model sizes. This suggests that combinatorial creativity requires a delicate balance between (1) models that are too shallow and wide, where insufficient depth may hinder the sequential processing capacity to handle in-memory leaps of thought (which are required to make distant, constrained associations between concepts) and (2) models that are too deep and narrow, which suffer from restricted representational capacity that may limit their ability to hold and associate the diverse concepts needed for novel combinations. Future work can use our framework as a starting point to explore this depth-width tradeoff in more detail mechanistically.
@@
In Figure 4, we plot the relationship between novelty and utility across all three model sizes. Previously, Varshney (2019) established a fundamental, information-theoretic limit between novelty and utility for combinatorial creativity. We find a similar novelty-utility tradeoff holds here: across all three scales, as the number of utility constraints increases, the novelty of artifacts exhibits a clear downward trend. While this tradeoff does not improve by increasing model size to 100M, frontier models today are well into the billions of parameters. Our work provides a foundation for future studies to explore this tradeoff for billion-parameter models.
@@
A series of recent studies have attempted to apply combinatorial creativity explicitly for scientific idea generation (Radensky et al., 2024, Sternlicht and Hope, 2025, Zhao et al., 2025). With the novelty-utility tradeoff in mind, we provide a potential explanation for why LLMs excel at generating novel research ideas (Si et al., 2024, Sanyal et al., 2025, Gu and Krenn, 2024, Wang et al., 2024, Guo et al., 2025) but struggle at ensuring their practical feasibility, in what has been termed the ideation-execution gap (Si et al., 2025). In Table 2, we explain how exclusion constraints can be viewed as a minimal abstraction for preventing unrealistic assumptions and excluding prohibitively expensive execution plans, while inclusion constraints can represent ensuring that a proper baseline is included and can serve as a minimal abstraction to ensure implementation plans are sufficiently detailed. Since the novelty-utility tradeoff remains persistent even at the 100M scale (see Figure 4), this suggests that the same fundamental tradeoff might plague the frontier models used in previous works, although a large-scale study pretraining at frontier-model scale should be performed to validate this explicitly. This finding is consistent with recent work from Shashidhar et al. (2025), which also identified a validity-diversity tradeoff in LLM-generated evaluation questions, where models that produced the most diverse (novel) questions often did so at the cost of lower factual validity (utility).